A particle start moving from rest state along a straight line under the action the action of a constant force and travel distance x in the first five seconds . The distance travelled by it in next five seconds will be.
 

Grade:11

1 Answers

Hasan Naqvi
8 years ago
We know that,
s= ut + 1/2 at2
u=0, and a= constant
 
x= ½ a * 5 * 5 = 12.5 a
Distance traveled in 10s = ½ * a * 10 * 10 = 50 a= 4x
 
So distance travelled between 5-10s = 3x